OpenStudy (anonymous):

If c=5/9(f-32),solve for f

OpenStudy (nenadmatematika):

do you mean c=5/(9*(f-32)) or 5*(f-32)/9

OpenStudy (nenadmatematika):

?

OpenStudy (anonymous):

c=5/9 (f-32) solve for f . That is how the proffesor has it written

OpenStudy (nenadmatematika):

hmmm OK f-32=9/5 c so f=9/5 c+32